OpenStudy (anonymous):

Find the volume of a rectangular prism with edges of 5d,6d, and 2d centimeters.

OpenStudy (anonymous):

\( 60 d^3 \)

OpenStudy (anonymous):

@FoolForMath o- o ..

OpenStudy (anonymous):

60d^3

OpenStudy (anonymous):

i was gonna say this ^ but she answer'd it Answer => 60d^3

OpenStudy (anonymous):

60d^3

OpenStudy (anonymous):

can't you guys see my answer?

OpenStudy (anonymous):

thanks guys!

OpenStudy (anonymous):

This is 2012!

OpenStudy (anonymous):

V = Area of rectangular * height = L * W * H = 5d * 6d * 2d = 50d³ cm³